A laser emits a narrow beam of light. The radius of the beam is 5.6 mm, and the power is 1.2 mW. What is the intensity of the laser beam?
Answer #1

Intensity = Power per unit Area

= Power / Area of radius of a circle
= P / pi*r^2.
= 1.2 x 10^-3 W / pi*(5.6x10^-3)^2
= 12.18 W/m^2
Answer #2

I = 12*10^-3/(pi*0.0056^2)

= 121.8 W/m^2
Answer #3

In physics, intensity is the power transferred per unit area


Intensity = 1.2/Area

Area = 4?r^2 = 393.88 * 10^-6


Intensity = 3.046 W/m^2
Answer #4

we know that ,

I = P/(4pr2)

= 1.2 * 10^-3 W / ( 4* 3.14* 0.0056m^2 )

= 3.04660081 W/m^2


therefore, intensity of laser beam is 3.04660081 W/m^2.